OpenStudy (anonymous):

S = p/3 + 7q solve for p

OpenStudy (anonymous):

3/ (s-7q)

OpenStudy (anonymous):

move the 7q over... (subtract 7q from both sides)

OpenStudy (anonymous):

multiply by 3... then take the inverse...

OpenStudy (anonymous):

so i was wrong in my first post... 1/ 3(s-7q) is the answer

jhonyy9 (jhonyy9):

S=p/3 +7q if just 3 is denominator so than you get S-7q=p/3 and from this if you multiply both sides by 3 you will get 3(S-7q)=p p=3(S-7q)=3S-21q